# A Scalable Matrix Visualization for Understanding Tree Ensemble Classifiers

### v1

@@ VI-B2 Quantitative Stock Trading
Overview. Initially, E3 wanted to verify whether the model utilized appropriate attributes for prediction. He checked the most frequently used attributes in the matrix view and found the top-5 of them are technical analysis attributes (Fig. 10(a)). E3 explained that these technical analysis attributes offered a more direct reflection of price fluctuations than fundamental attributes. So, the model relied more on these technical analysis attributes to make predictions.
Next, E3 investigated whether the key attributes were correctly used. He examined the most important attribute KLEN in the first column (Fig. 10A), which reflects the previous day’s stock price range. He quickly found that the model tended to predict the stock price as “stable” when KLEN was low. This is reasonable because stocks with lower KLEN are more likely to maintain stable prices in the next few days.
@@
Addressing biased predictions. When analyzing the attribute KLEN, E3 identified an exception: a rule made an “increase” prediction when KLEN was low (Fig. 10B). In addition to KLEN, this rule used three other attributes: RESI20, industry, and RESI60. Examining the samples covered by this rule, he found that these samples had small KLEN, medium RESI20, and medium RESI60, all of which indicated stable stock prices. However, the condition of industry=Bank may be the reason for the prediction of “increase.” E3 commented that the economic cycles of different industries can vary a lot and lead to biased predictions. Specifically, the banks associated with these samples experienced a growth period when the training data was collected, so the trained model tended to predict them as “increase.” However, such a pattern did not exist in the test samples and thus caused a performance drop. To address this issue, he removed the attribute industry. After retraining the model, he observed an improvement in accuracy from 56.26% to 56.52%. Specifically, on the samples with industry=Bank, the accuracy increased from 48.90% to 51.22%.
@@
Improving model performance in “increase.” As E3’s primary goal was to find stocks with possibly rising prices to maximize profits, he focused on analyzing the attributes related to the “increase” class in the retrained model. As shown in Fig. 10C, he observed that STD60 was frequently used for the prediction of “increase”. After investigating the rules using this attribute, he noticed that the rules with the conditions of larger STD60 consistently made predictions of “increase” (Fig. 10D). However, E3 commented that a larger STD60 just meant a large fluctuation in stock prices over the past 60 days, which did not necessarily indicate an increasing stock price. Given the heavy reliance on this attribute to predict “increase” and the low prediction accuracy in this class (26.35%), E3 concluded that STD60 was not a useful attribute for the prediction of this class and more attributes were needed. He then performed feature engineering by leveraging TA-lib [3], an open-source tool, to generate more technical analysis attributes. After adding these attributes and retraining the model, the overall accuracy increased from 56.52% to 58.35%, and the class accuracy of “increase” increased from 26.35% to 28.24%.
@@
To verify whether the newly added attributes were used correctly, E3 examined the extracted rules again. The usage of STD60 decreased from 26 to 7 (Fig. 10E), while a newly added attribute, NATR, became the most frequently used (Fig. 10F). This attribute represents asset volatility. A high NATR indicates a large variation in stock prices. Sorting rules by NATR, E3 found that, indeed, high NATR values are associated with “increase” or “decrease” predictions, while low NATR values are associated with “stable” predictions. (Fig. 10G) E3 wondered how other attributes collaborated with high NATR to predict the “increase” or “decrease.” Therefore, he explored rules with the condition of high NATR. The neighborhood was dynamically constructed, where he found that revenueGrowth, the growth of the profit rate, was frequently used in these rules (Fig. 10H). The ordering result reveals a clear pattern that high revenueGrowth mostly led to an “increase” prediction, while low revenueGrowth mostly led to a “decrease” prediction. This observation was consistent with his experience.
@@
Removing redundant attributes. During further exploration, E3 identified some redundant attributes for prediction. For example, VMA20, which indicates the average trading volume in the preceding 20 days, was recognized as an important attribute and ranked the ninth among the over 100 attributes (Fig. 10I). However, in the rules using this attribute, both low and high VMA20 resulted in the predictions of “stable” (Fig. 10J), which seemed to suggest the value of VMA20 does not affect the prediction. To further analyze these rules, E3 pinned VMA20 to the first column and sorted rules based on it. He found that all the rules using VMA20 also used NATR and had low NATR values. As low NATR indicates a small variation in stock prices, E3 believed it was the actual reason behind the predictions of these rules, and VMA20 was redundant. He proposed to remove VMA20. Similarly, he removed another eight redundant attributes. After retraining the model, its accuracy was improved to 59.20%.
@@ VI-C Expert Feedback and Discussion
After the case studies, we conducted semi-structured interviews with each of the six experts (E1-E6) introduced in Sec. III-A. As E4-E6 were not involved in the case studies, we first introduced RuleExplorer and presented the case studies to them. All the experts were required to use the tool to understand and verify models, and identify problems. We also discussed the advantages and limitations of the tool. Each interview lasted about 35–60 minutes.
@@ VI-C1 Usability
Learning curve. The experts consider RuleExplorer is easy to use. They like that the matrix view takes a format similar to a data table and naturally supports table operations, such as sorting and swapping. This familiar design has enhanced their analysis efficiency. They also indicate that using RuleExplorer enables them to quickly understand the model, which has increased their trust in the model and their confidence in decision-making. Specifically, they expressed satisfaction with the dynamic hierarchy, which enables them to analyze different subsets of rules on demand. The experts also value the ease of finding anomalous rules using RuleExplorer. As E2 commented, “Different from RuleMatrix [34] and ExplainableMatrix [37], this tool automatically recommends some unusual rules, which help me quickly identify potential issues in the model and make interventions accordingly.”
Scalability. Matrix visualizations often face the scalability issue due to the limited number of rows and columns that can be displayed [33, 58]. To tackle this issue, RuleExplorer 1) organizes a large rule set into a hierarchy and dynamically selects representative rules to display, and 2) sorts attributes by importance and divides them into pages. These two strategies are effective because users usually focus on a small subset of rules and key attributes. In addition, our tool can easily be extended to handle other scalability issues. For example, for attributes with thousands of categories, we can employ alternative feature transformation methods, such as Target Encoding [19], which replaces each category with a single numerical value representing its statistical correlation to the target variable. For classification tasks involving more than 10 classes, we can hierarchically organize classes and dynamically assign colors during user exploration  [10].

### v2

@@ VI-B2 Quantitative Stock Trading
Overview. Initially, E3 wanted to verify whether the model utilized appropriate attributes for prediction. He checked the most frequently used attributes in the matrix view and found the top-5 important attributes are technical analysis attributes (Fig. 10(a)). E3 considered it reasonable as technical analysis attributes offered a more direct reflection of price fluctuations than fundamental attributes. Reordering rules based on the most important attribute KLEN, he found that the model tended to predict the stock price as “stable” when KLEN was low. This is reasonable because stocks with lower KLEN are more likely to maintain stable prices in the next few days.
@@
Addressing biased predictions. However, there was an exception: when KLEN was low, one rule made an “increase” prediction (Fig. 10B). Examining the samples covered by this rule, E3 found that these samples had small KLEN, medium RESI20, and medium RESI60, all of which indicated stable stock prices. However, they also shared the condition of industry=Bank, which may be the reason for the prediction of “increase.” E3 explained that the banks associated with these samples experienced a growth period when the training data was collected, so the trained model tended to predict them as “increase.” However, such a pattern did not exist in the test samples and thus caused a performance drop. To address this issue, he removed the attribute industry. After retraining the model, he observed an improvement in accuracy from 56.26% to 56.52%. Specifically, on the samples with industry=Bank, the accuracy increased from 48.90% to 51.22%.
@@
Improving model performance in “increase.” As E3’s primary goal was to find stocks with possibly rising prices to maximize profits, he focused on analyzing the attributes associated to the “increase” prediction in the retrained model. As shown in Fig. 10C, he observed that STD60 was frequently used for the prediction of “increase”. Reordering rules based on this attribute, he noticed that the rules with the conditions of larger STD60 consistently made predictions of “increase” (Fig. 10D). However, as E3 commented, a larger STD60 just meant a large fluctuation in stock prices over the past 60 days, which did not necessarily indicate an increasing stock price. E3 considered STD60 was not a useful attribute for the “increase” prediction, and suspected the heavy reliance on it accounted for the low prediction accuracy in this class (26.35%). He then performed feature engineering by leveraging TA-lib [4], an open-source tool, to generate more technical analysis attributes. After adding these attributes and retraining the model, the overall accuracy increased from 56.52% to 58.35%, and the class accuracy of “increase” increased from 26.35% to 28.24%.
@@
To verify whether the newly added attributes were used correctly, E3 examined the cumulative statistics for attributes at the bottom of the matrix view. The usage of STD60 decreased from 26 to 7 (Fig. 10E), while a newly added attribute, NATR, became the most frequently used (Fig. 10F). NATR represents asset volatility. A high NATR indicates a large variation in stock prices. Reordering rules based on NATR, E3 found that, indeed, high NATR values are associated with “increase” or “decrease” predictions, while low NATR values are associated with “stable” predictions (Fig. 10G). Examining the neighborhood of rules with high NATR, E3 found that revenueGrowth, the growth of the profit rate, was frequently used together (Fig. 10H). Reodering rules based on NATR and revenueGrowth revealed a clear pattern that high revenueGrowth mostly led to an “increase” prediction, while low revenueGrowth mostly led to a “decrease” prediction. This observation was consistent with his experience.
@@
Removing redundant attributes. During further exploration, E3 identified some redundant attributes for prediction. For example, VMA20, which indicates the average trading volume in the preceding 20 days, was recognized as an important attribute (Fig. 10I). However, rules using this attribute made predictions of “stable” no matter low or high VMA20 values, which seemed to suggest the value of VMA20 does not affect the prediction. Pinning VMA20 to the first column and reordering rules based on it, E3 found that all the rules using VMA20 also had low NATR values. As low NATR indicates a small variation in stock prices, E3 believed it was the actual reason for the predictions of these rules, and VMA20 was redundant. He proposed to remove VMA20. Similarly, he removed another eight redundant attributes. After retraining the model, its accuracy was improved to 59.20%.
@@ VI-C Expert Feedback and Discussion
After the case studies, we conducted semi-structured interviews with each of the six experts (E1-E6) introduced in Sec. III-A. E4-E6 were not involved in the case studies. These newly involved experts included a fund manager with two years of experience (E4), a vice president at a commercial bank branch (E5), and a machine learning researcher with five years of experience (E6). For E4-E6, we spent the initial 20 minutes introducing RuleExplorer and presenting the case studies. Each expert then used RuleExplorer for model understanding and verification tasks. It took 5-15 minutes for them to become familiar with the interactions and conduct analysis. We concluded with a discussion on RuleExplorer’s advantages and limitations. Each interview lasted 35-60 minutes.
@@ VI-C1 Usability
Learning curve. The experts consider RuleExplorer easy to use. They like that the matrix view takes a format similar to a data table and naturally supports table operations, such as sorting and swapping. This familiar design has enhanced their analysis efficiency. They also indicate that using RuleExplorer enables them to quickly understand the model, which has increased their trust in the model and their confidence in decision-making. Specifically, they expressed satisfaction with the dynamic hierarchy, which enables them to analyze different subsets of rules on demand. The experts also value the ease of finding anomalous rules using RuleExplorer. As E2 commented, “Different from RuleMatrix [40] and ExplainableMatrix [43], this tool automatically recommends some unusual rules, which help me quickly identify potential issues in the model and make interventions accordingly.” We believe that conducting a user study to compare RuleExplorer with these interactive systems would provide deeper insights into its usability and its effectiveness in facilitating model understanding. We consider this an important future work.
Scalability. Matrix visualizations often face the scalability issue due to the limited number of rows and columns that can be displayed [39, 65]. To tackle this issue, RuleExplorer 1) organizes a large rule set into a hierarchy and dynamically selects representative rules to display, and 2) sorts attributes by importance and divides them into pages. These two strategies are effective because tree ensemble models typically rely on a smaller set of key attributes for decision making [7], and users usually focus on a small subset of representative rules and key attributes for initial analysis. In addition, our tool can easily be extended to handle other scalability issues. For example, for attributes with thousands of categories, we can employ alternative feature transformation methods, such as Target Encoding [24], which replaces each category with a single numerical value representing its statistical correlation to the target variable. For classification tasks involving more than 10 classes, we can hierarchically organize classes and dynamically assign colors during user exploration  [15].